# BABILong: Testing the Limits of LLMs with Long Context Reasoning-in-a-Haystack

### v1

@@ 4 Related Work on Long Context Benchmarks and Datasets
To further extend benchmarks lengths with real and human annotated data is very challenging. Therefore, "needle-in-a-haystack" inspired benchmarks arised. Following LLMTest  [^5] https://github.com/gkamradt/LLMTest_NeedleInAHaystack with magic numbers as needles in Paul Graham essays as haystack, passkey and key-value retrieval tasks are part of InfinityBench (Zhang et al., 2024b). Counting-Stars (Song et al., 2024b) suggests to insert multiple sentences about little penguins that count stars into the same essays for English or The Story of the Stone for the Chinese language. The task is to answer questions based on these "needle" sentences. RULER (Hsieh et al., 2024) extends "needle-in-a-haystack" with multiple types and amount of "needles". RULER and Counting-Start introduce new task categories such as multi-hop tracing and aggregation to test models beyond searching from context.
@@
BABILong focuses on natural language reasoning over multiple facts distributed in very large textual corpora. Compared to existing approaches it provides more tasks and more natural and deceptive mixing of information into background documents. BABILong consists of diverse set of 20 tasks that cover different capabilities including multi-hop tracing, aggregation over needles and extending them with basic deduction and induction, time, positional, and size reasoning, and path finding. The benchmark goes with predefined splits up to unprecedented 1M token length. Lengths beyond 1M tokens could be generated and we test models up to 11M tokens. Furthermore, while we evaluate models on English-only tasks from bAbI (Weston et al., 2016) adding new languages is straightforward.
@@ Conclusions
In this work, we introduced the BABILong, a diverse and scalable benchmark designed to bridge the gap in evaluating large language models (LLMs) across extensive context lengths. Our experiments demonstrate that BABILong offers a more representative evaluation framework for long-context reasoning among the existing benchmarks. The analysis of correlation with other benchmarks further validates BABILong’s ability to pose a significant challenge for large language models to maintain performance as context lengths scale up. The BABILong benchmark offers algorithmically adaptable document length and facts placement, includes predefined sets of bins ranging from 0k to 1M tokens. Facts in BABILong could be generated making it leak-proof for future LLMs. It consists of a set of 20 diverse tasks covering reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets. Compared to other benchmarks, BABILong shows high correlation on short contexts with MMLU and diverges from it as lengths increases.
Our findings reveal limitations of popular open-source LLMs as well as GPT-4 and RAG regarding effective long context utilization. Their performance heavily relies on the first 5-25 % of the input, highlighting the need for improved context processing mechanisms. Fine-tuning experiments show that tasks from BABILong are solvable even by relatively small models like RMT with GPT-2 (137M) and Mamba (130M). Fine-tuning improves the performance of GPT-3.5-Turbo and Mistral-7B, but their context lengths remain limited to 16K and 32K, respectively. Among the evaluated models, Mamba (130M) and RMT(137M) achieve the strongest results. However, Mamba is hard to infer on lengths more than 128K tokens, while RMT enables the processing of lengths up to 11 million tokens.
@@ Appendix B Related Work on Long Context Models
Appendix C Details on RMT and Mamba fine-tuning on BABILong
@@
We used the GPT-2 (Radford et al., 2019) (137M) model as the backbone for RMT. The segment size was fixed at 512 tokens, and the model was augmented with 16 memory tokens. Finetuning was conducted on BABILong using a curriculum schedule with progressively increasing sequence lengths: 1, 2, 4, 6, 8, 16 and 32 segments. For each curriculum step $n$ we chose the number of segments randomly from 1 to $n$ for every batch to prevent overfitting to a certain context size. We maintained a fixed batch size of 64 and the AdamW Loshchilov & Hutter (2019) optimizer with learning rate 1e-05, a linear schedule and 1000 warmup steps. Each curriculum stage had a maximum of 10,000 steps, although checkpoints were frequently finalized earlier based on the best validation exact match metric. The weight decay value was set to 0.01, and no gradient truncation was used. Training was performed on 1-4 Nvidia A100 or H100 GPUs with the duration of each curriculum stage ranging from 40 minutes to 20 hours.
@@
Appendix D Detailed LLM evaluation on BABILong tasks
Here we present the complete results of LLM evaluation. Table 3 showcases the performance of 27 models across the first five tasks. Comparing the tasks in the table makes evident the difference in task complexity for language models. QA1 and QA5 are the easiest, with most models achieving over 70% accuracy for the 0k split. QA4 is significantly more challenging, and only 5 models can reach this level of performance. QA2 and QA3 pose even greater challenges for most models.
The number of parameters positively impacts accuracy on the shortest 0k split. Among not-finetuned models, GPT-4, Phi-3-medium, Jamba, Command-R, Yi-34B and Mixtral 8x22B consistently outperform smaller models. Notably, RWKV and Mamba-2.8B also demonstrate strong performance on QA2 and QA3. However, as the context length increases, some of the largest models lose their advantage over smaller ones.
It is important to note, that all BABILong tasks are in practice solvable even with smaller models. Finetuned RMT and Mamba achieve outstanding scores across most sequence lengths, significantly outperforming LLMs despite having up to 100 times ewer parameters. Mamba has an advantage on medium-length sequences, but RMT excells in processing much larger sequences up to 10M tokens.
Appendix E RMT Training and Evaluation Details
For RMT we use curriculum training with sequentially increasing number of segments. RMT uses the following schedule for number of segments: 1-2-4-6-8-16-32. During each curriculum stage $n$ the number of segment is chosen randomly from $1$ to $n$. We select learning rate from {5e-05, 3e-05} and use the AdamW optimizer and linear learning rate scheduling with warmup. We use a total batch size of $64$ and train for {5000, 10000} steps with early stopping if metrics stop increasing. For the backbone transformer we use the pretrained GPT-2 137M from HuggingFace: https://huggingface.co/GPT-2. We used up to 4 Nvidia A100 80Gb per experiment.
We evaluate models on the BABILong benchmark. We use full test set for sequences up to 1M tokens, for 10M tokens we only provide results for 100 samples. As shown in Table 4, evaluation time grows linearly with context length. We fix a random seed used to sample background texts from PG19 for the test set. However, the seed was not fixed to avoid overfitting to specific sampled texts during training.
@@ Appendix H Recurrent Memory Transformer Analysis
Appendix I LLM fine-tuning
@@
Appendix J Prompts Used to Benchmark GPT-4-Turbo and Mistral Models
We used the same prompts to evaluate GPT-4-Turbo and Mistral models in our tasks. Each prompt starts with the description of the task followed by several examples inside the $<$example$>$ $<$/example$>$ tags. The next section inside $<$context$>$ $<$/context$>$ tags contains an instance of the task. We additionally duplicate the question with the QUESTION mark, in order for the model recognize the question in the large input prompts. The last sentences specify the required response format.
@@
The BABILong benchmark is designed to test language models’ ability to reason across facts distributed in extremely long documents. BABILong includes a diverse set of 20 reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets. Today, Large language models (LLMs) and neural architectures are continually evolving and achieving remarkable improvements, particularly in their ability to handle longer contexts, but the benchmarks used to evaluate them have not kept pace. For example, current benchmarks such as Longbench (Bai et al., 2023) and L-Eval An et al. (2023) scale only up to 40,000 tokens, while models are capable of hundreds of thousands and millions of tokens (OpenAI, 2023; Bulatov et al., 2024; Gu & Dao, 2023; Anthropic, 2024; Reid et al., 2024; Liu et al., 2024a). To bridge this gap, BABILong allows the construction of tasks of almost arbitrary length, in order to adapt them to the evaluation of new, more powerful models in an extensible and controllable way.
@@
This information would be added in camera-ready version of the paper.

### v2

@@ 4 Related Work on Long Context Benchmarks and Datasets
Further extending of benchmarks’ length with real and human annotated data is very challenging. Therefore, "needle-in-a-haystack" inspired benchmarks were proposed. Following LLMTest  [^6] https://github.com/gkamradt/LLMTest_NeedleInAHaystack with magic numbers as needles in Paul Graham essays as haystack, passkey and key-value retrieval tasks are part of InfinityBench (Zhang et al., 2024b). Counting-Stars (Song et al., 2024b) suggests to insert multiple sentences about little penguins that count stars into the same essays for English or The Story of the Stone for the Chinese language. The task is to answer questions based on these "needle" sentences. RULER (Hsieh et al., 2024) extends "needle-in-a-haystack" with multiple types and amount of "needles". RULER and Counting-Stars introduce new task categories such as multi-hop tracing and aggregation to test models beyond searching from context.
@@
BABILong focuses on natural language reasoning over multiple facts distributed in very large textual corpora. Compared to existing approaches it provides more tasks and more natural and deceptive mixing of information into background documents. BABILong consists of diverse set of 20 tasks that cover different capabilities including multi-hop tracing, aggregation over needles and extending them with basic deduction and induction, time, positional, and size reasoning, and path finding. The benchmark goes with predefined splits up to unprecedented 10M token length. Lengths beyond 10M tokens could be generated and we test models up to 50M tokens. Furthermore, while we evaluate models on English-only tasks from bAbI (Weston et al., 2016) adding new languages is straightforward.
@@ Conclusions
In this work, we introduced the BABILong, a diverse and scalable benchmark designed to bridge the gap in evaluating large language models (LLMs) across extensive context lengths. Our experiments demonstrate that BABILong offers a more representative evaluation framework for long-context reasoning among the existing benchmarks. The analysis of correlation with other benchmarks further validates BABILong’s ability to pose a significant challenge for large language models to maintain performance as context lengths scale up. The BABILong benchmark offers algorithmically adaptable document length and facts placement, includes predefined sets of bins ranging from 0k to 10M tokens. Facts in BABILong could be generated making it leak-proof for future LLMs. It consists of a set of 20 diverse tasks covering reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets. Compared to other benchmarks, BABILong shows high correlation on short contexts with MMLU and diverges from it as lengths increases.
Our findings reveal limitations of popular open-source LLMs as well as GPT-4, Gemini 1.5 Pro and RAG solutions regarding effective long context utilization. Their performance heavily relies on the first 5-25% of the input, highlighting the need for improved context processing mechanisms. The recent open-source models LLama-3.1 and Qwen-2.5 show the best performance of pre-trained LLMs, with LLama-3.1 using pre-training and supervised fine-tuning on long context data up to 128K tokens. BABILong fine-tuning experiments show that tasks from BABILong are solvable even by relatively small models like RMT & ARMT with GPT-2 (137M) and Mamba (130M). Fine-tuning improves the performance of GPT-3.5-Turbo and Mistral-7B, but their context lengths remain limited to 16K and 32K, respectively. Among the evaluated models, Mamba and recurrent transformers achieve the strongest results. However, Mamba is hard to infer on lengths more than 128K tokens, while RMT and ARMT enables the processing of lengths up to 50 million tokens.
@@
We could also use other sources of facts and questions (e.g., reasoning datasets such as Ruletaker (Clark et al., 2020), ProofWriter (Tafjord et al., 2021), FOLIO Han et al. (2022), PrOntoQA (Saparov & He, 2023), LogicNLI (Tian et al., 2021), and DELTA (Poulis et al., 2024)), mixing samples of question-answering datasets with background text from the same domain, or using LLMs to generate questions about statements that belong to the original documents. Keeping the same framework as in BABILong, this will lead to more complex and real-world scenarios.
@@
Author contributions
Y.K., A.B. and M.B. conceived of the idea. A.B. prepared the dataset. Y.K. and A.B. scored open-source LLMs, A.B. scored RMT, I.R. scored ARMT, Mamba and Gemini pro 1.5, P.A. scored Mistal models and RAG pipelines, D.S. scored GPT-4 and GPT-3.5, A.S. fine-tuned and scored GPT-3.5, A.B. finetuned and scored Mistral. D.S. created public leaderboard. A.B. populated and maintain leaderboard. Y.K. performed study to compare MMLU, RULER and BABILong. Y.K., A.B. and M.B. aggregated and analyzed scoring results. All authors contributed to discussions of results. Y.K., A.B. and P.A. wrote the first draft of the manuscript. All the authors contributed in extending manuscript with their results. M.B. took the lead in structuring and editing manuscript towards the final version.
Acknowledgments and Disclosure of Funding
We are thankful to SberDevices for granting us access to additional computational resources. This work was partially supported by a grant for research centers, provided by the Analytical Center for the Government of the Russian Federation in accordance with the subsidy agreement (agreement identifier  [^00] D730324P540002) and the agreement with the Moscow Institute of Physics and Technology dated November 1, 2021 No. 70-2021-00138.
Appendix Content
Appendix A Code and Data Availability
Appendix B Related Work on Long Context Models
Appendix C Details on RMT and Mamba fine-tuning on BABILong
Appendix D Detailed LLMs evaluation on BABILong QA1-5 tasks
Appendix E Gemini Evaluation
Appendix F BABILong Dataset Statistics
Appendix G Details of the RAG Pipeline
Appendix H Recurrent Memory Transformer Analysis
Appendix I LLMs fine-tuning results
Appendix J Prompts Used to Benchmark Large Language Models
Appendix L BABILong Task Examples
Appendix M Author Statement
Appendix N BABILong Datasheet
@@ Appendix B Related Work on Long Context Models
Appendix C Details on RMT, ARMT, and Mamba fine-tuning and evaluation on BABILong
@@
We used the GPT-2 (Radford et al., 2019) (137M) model (https://huggingface.co/GPT-2) as the backbone for RMT and ARMT. The segment size was fixed at 512 tokens, and the model was augmented with 16 memory tokens for RMT and 10 memory tokens in ARMT. Finetuning was conducted on BABILong using a curriculum schedule with progressively increasing sequence lengths: 1, 2, 4, 6, 8, 16 and 32 segments. ARMT used 2-3-5-8-16-32. For each curriculum step $n$ we chose the number of segments randomly from 1 to $n$ for every batch to prevent overfitting to a certain context size. We maintained a fixed batch size of 64 and the AdamW Loshchilov & Hutter (2019) optimizer with learning rate in range {5e-05, 3e-05}, a linear schedule and 1000 warmup steps. In ARMT experiments, the learning rate was set to 1e-04. We also consider the memory-dimension parameter for ARMT to be 64 and we use non-linearity DPFP-3 (Schlag et al., 2021). Each curriculum stage had a maximum of 10,000 steps with early stopping if metrics stop increasing. The weight decay value was set to 0.01, and no gradient stopping was used. Training was performed on 1-4 Nvidia A100 or H100 GPUs with the duration of each curriculum stage ranging from 40 minutes to 20 hours.
@@
To evaluate recurrent models on the BABILong benchmark we use the full test set for sequences up to 1M tokens, for 10M tokens we only provide results on averaged over 100 samples. As shown in Table 3, evaluation time grows linearly with context length. We fix a random seed used to sample background texts from PG19 for the test set. However, the seed was not fixed to avoid overfitting to specific sampled texts during training.
@@
Appendix D Detailed LLMs evaluation on BABILong QA1-5 tasks
Here we present the complete results of LLMs evaluation. Table 4 showcases the performance of 38 models across the first five tasks. Comparing the tasks in the table makes evident the difference in task complexity for language models. QA1 and QA5 are the easiest, with most models achieving over 70% accuracy for the 0k split. QA4 is significantly more challenging, and only 5 models can reach this level of performance. QA2 and QA3 pose even greater challenges for most models.
The number of parameters positively impacts accuracy on the shortest 0k split. Among not-finetuned models, GPT-4, Phi-3-medium, Qwen, Jamba, Command-R, Yi-34B and Mixtral 8x22B consistently outperform smaller models. Notably, RWKV and Mamba-2.8B also demonstrate strong performance on QA2 and QA3. However, as the context length increases, some of the largest models lose their advantage over smaller ones.
It is important to note, that all BABILong tasks are in practice solvable even with smaller models. Finetuned RMT, ARMT, and Mamba achieve outstanding scores across most sequence lengths, significantly outperforming LLMs despite having up to 100 times ewer parameters. Mamba has an advantage on medium-length sequences, but recurrent memory models (RMT and ARMT) excel in processing much larger sequences up to 10M tokens.
Appendix E Gemini Evaluation
Figure 6: Results of Gemini 1.5 Pro 002 evaluations. Built-in content safety filtering causes up to 14% of requests to be denied with increased context size. Without rejected means that we removed from evaluation all cases where the model refused to answer. Few-shot means that we used instructions with in-context examples.
We evaluated the Gemini 1.5 Pro 002 model on the QA1 task of BABILong. We were requesting model responses via API, some of the requests were denied due to built-in content safety filtering, even with BLOCK_NONE set. In Fig. 1b we only report results for requests where the model did not refused to response. We present the full results in Figure 6. We found that as the context size increases, the model can refuse to respond up to 14% of the time. We used 1000 samples for lengths up to 32K, for larger lengths we used 100 samples per length.
@@ Appendix H Recurrent Memory Transformer Analysis
Appendix I LLMs fine-tuning results
@@
Appendix J Prompts Used to Benchmark Large Language Models
Each prompt starts with the description of the task followed by several examples inside the $<$example$>$ $<$/example$>$ tags. The next section inside $<$context$>$ $<$/context$>$ tags contains an instance of the task. We additionally duplicate the question with the QUESTION mark, in order for the model recognize the question in the large input prompts. The last sentences specify the required response format.
@@
The BABILong benchmark is designed to test language models’ ability to reason across facts distributed in extremely long documents. BABILong includes a diverse set of 20 reasoning tasks, including fact chaining, simple induction, deduction, counting, and handling lists/sets. Today, Large language models (LLMs) and neural architectures are continually evolving and achieving remarkable improvements, particularly in their ability to handle longer contexts, but the benchmarks used to evaluate them have not kept pace. For example, current benchmarks such as Longbench (Bai et al., 2023) and L-Eval An et al. (2023) scale only up to 40,000 tokens, while models are capable of hundreds of thousands and millions of tokens (OpenAI, 2023b; Bulatov et al., 2024; Gu & Dao, 2023; Anthropic, 2024; Reid et al., 2024; Liu et al., 2024a). To bridge this gap, BABILong allows the construction of tasks of almost arbitrary length, in order to adapt them to the evaluation of new, more powerful models in an extensible and controllable way.
@@
Refer to the acknowledgments section of the main text.